# Association and Independence Test for Random Objects

### v2

@@ 3.1 Testing independence for random objects
We are interested in testing the hypothesis
[…]
using random samples $\mathcal{D}_{\mathsf{XY}}:=\{(X_{i},Y_{i}):1\leq i\leq n\}$ drawn from the joint distribution of the random objects $(X,Y)$. The profile association $\Delta(X,Y)$ in (4) quantifies the strength of the relationship between random objects $X$ and $Y$ and provides the foundation for a test statistic to assess their independence.
Consider the functions $\psi_{1}:\mathcal{M}_{1}^{\otimes 4}\mapsto\{-1,0,1\}$ defined as
@@
×$\quad\times\psi(y_{1},y_{2},y_{3},y_{4})\psi(y_{1},y_{2},y_{5},y_{6}).$
In what follows, we use the notation $\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{6},Y_{6})\}$ to denote six independent copies of $(X,Y)$, where one has
@@
Given random samples $\mathcal{D}_{\mathsf{XY}}$, we employ a U-statistic as an estimator of $\Delta(X,Y)$, defined as
$D_{n}(\mathcal{D}_{\mathsf{XY}})=\frac{(n-6)!}{n!}\sum_{\mathcal{P}_{6}^{n}(\bm{\alpha})}h\left((X_{\alpha_{1}},Y_{\alpha_{1}}),\ldots,(X_{\alpha_{6}},Y_{\alpha_{6}})\right),$ (5)
where $\sum_{\mathcal{P}_{6}^{n}(\bm{\alpha})}$ demotes summation over all $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{6})^{\top}$ such that
$\alpha_{i}=1,\ldots,n;\,\,\alpha_{i}\neq\alpha_{j}\text{ if }i\neq j,\,\,\text{ for all }\,i,j=1,2,\ldots,6.$
We require some concepts related to the analysis of U-statistics, which were first introduced by Hoeffding (1948), with further details provided in Serfling (2009). Since the function $h$ is not symmetric, we first symmetrize it as follows,
@@
where $\mathcal{P}_{6}^{6}(\bm{\alpha})$ denotes the set of all permutations of indices $\{1,2,\ldots,6\}$. For $r=1,2,\ldots,6$, let
@@
[…]
@@
for all $(u,v)\in\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ and $f\in\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$.
@@ 3.2 Permutation test and asymptotic results
Since the asymptotic distribution in Theorem 2 is unknown, we use a permutation test to determine the critical values of the test statistic $D_{n}(\mathcal{D}_{\mathsf{XY}})$ when implementing the test. Permuting the data and recalculating the test statistic for each permutation is a standard approach to generate a reference distribution. In testing independence, the typical approach is to compute replicates of the test statistic by randomly permuting the $Y$ sample while keeping the order of the $X$ sample fixed (Székely et al., 2007). However, randomly permuting the entire $Y$ sample introduces dependence among the data pairs under the alternative $\mathcal{H}_{1}$ when $X$ and $Y$ are dependent, making it challenging to establish consistency of the critical value under $\mathcal{H}_{1}$. Here we adopt a half-permutation procedure introduced in Fan et al. (2024) that addresses this issue and does not require the splitting technique with its possible loss of power. Adopting this approach for our setting requires a different proof technique to obtain consistency and power bounds for the proposed profile independence test, as previous techniques are not applicable.
@@
We use $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$ to denote the set of permutations of $\{1,2,\ldots,n\}$ with cardinality $\lfloor n/2\rfloor$. Consider a sequence of i.i.d. random permutations $\Pi_{1},\ldots,\Pi_{B}$ that are independent of $(X,Y)$ and uniformly distributed on $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$. We use $\Pi_{j}^{c}$ to denote the complement of $\Pi_{j}$ with respect to the set $\{1,2,\ldots,n\}$. Without loss of generality, we assume $n$ is even, so that $\Pi_{j}$ and $\Pi_{j}^{c}$ have the same cardinality $n/2$. For each $\Pi_{j}^{c}$, let $\{\sigma_{jk}\}_{k=1}^{N_{j}}$ be a sequence of random permutations on $\Pi_{j}^{c}$ that are independent of $(X,Y)$ and $\Pi_{j}$. Assume $\Pi_{j}=\{\imath_{1},\imath_{2},\ldots,\imath_{\lfloor n/2\rfloor}\}$ and $\sigma_{jk}\circ\Pi_{j}^{c}=\{\jmath_{1},\jmath_{2},\ldots,\jmath_{\lfloor n/2\rfloor}\}$, where $\sigma_{jk}\circ\Pi_{j}^{c}$ denotes the sequence of $\Pi_{j}^{c}$ under the permutation $\sigma_{jk}$. The permuted data corresponding to $(\Pi_{j},\sigma_{jk})$ are
@@
$\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}):=\{(X_{\imath_{1}},Y_{\jmath_{1}}),(X_{\imath_{2}},Y_{\jmath_{2}}),\ldots,(X_{\imath_{\lfloor n/2\rfloor}},Y_{\jmath_{\lfloor n/2\rfloor}})\}.$
@@
The test statistic based on permuted data $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ is
@@
$D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})):=\frac{(\lfloor n/2\rfloor-6)!}{\lfloor n/2\rfloor!}\sum_{\mathcal{P}_{6}^{\lfloor n/2\rfloor}(\bm{\alpha})}h((X_{\imath_{\alpha_{1}}},Y_{\jmath_{\alpha_{1}}}),\ldots,(X_{\imath_{\alpha_{6}}},Y_{\jmath_{\alpha_{6}}})).$
@@
$\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}}):=\mathbb{P}\left(1+\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\{n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})\leq\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}}\leq(1+N)\alpha\right),$ (8)
where the probability in (8) is taken over $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and the only randomness in $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ comes from the random permutations $(\Pi_{j},\sigma_{jk})$.
Under the null hypothesis, $X$ and $Y$ are independent, and the elements in the data sets $\mathcal{D}_{\mathsf{XY}}$ and $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ are exchangeable. Consequently, $\{\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}_{j,k}$ and $n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})$ have the same asymptotic distribution by Theorem 2. Therefore, the probability on the right-hand side of equation (8) equals $\alpha$ for all distributions of $(X,Y)$ under the null hypothesis, which implies that $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ is a valid size-$\alpha$ test. For the consistency under the alternative $\mathcal{H}_{1}$, denote by $H(z)$ the cumulative distribution function of the asymptotic null distribution in Theorem 2. We employ the estimator
@@
for $H(z)$ based on the permuted data, where $N=\sum_{j=1}^{B}N_{j}$ is the total number of permutation pairs $(\Pi_{j},\sigma_{jk})$. Theorem 3 below establishes the consistency of this estimator, which ensures that the size of the proposed profile independence test is correctly controlled under the alternative.
@@ 4 Conditional independence test
Conditional independence testing is a statistical tool used to determine whether two variables are independent when conditioning on a third variable. It is used in a wide range of statistical applications where one studies complex relationships among variables, including causal inference (Zhang et al., 2011; Zhang et al., 2017; Peters et al., 2017), graphical models (Koller and Friedman, 2009; Williams and Mulder, 2020) and sufficient dimension reduction (Cook, 2004; Chen et al., 2015; Li, 2018). Many nonparametric conditional independence tests have been studied based on conditional density functions (Huang, 2010; Su and White, 2008; Banerjee et al., 2024), distribution functions (Linton and Gozalo, 1996; Zhou et al., 2020) and characteristic functions (Su and White, 2007; Wang et al., 2015; Wang and Hong, 2018). These methods are all restricted to the case where data reside in a vector space. To our knowledge, conditional independence tests for random elements in general metric spaces that do not support linear operations have remained unexplored.
In this section, we consider a setting where $X\in\mathcal{M}_{1}$ and $Y\in\mathcal{M}_{2}$ are random objects, while the conditional variable $Z$ is Euclidean, specifically $Z\in\mathcal{Z}$, where $\mathcal{Z}$ is a compact subset of ${\mathbb{R}}^{p}$. Our goal is to test the following hypothesis:
[…]
@@
$F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v):=\mathbb{P}_{X,Y,Z}\left(d(X,X^{\prime})\leq u,\,d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right),$
$F_{X^{\prime},Z^{\prime}}^{X}(u):=\mathbb{P}_{X,Z}\left(d(X,X^{\prime})\leq u\mid Z=Z^{\prime}\right),\,F_{Y^{\prime},Z^{\prime}}^{Y}(v):=\mathbb{P}_{Y,Z}\left(d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right).$
@@
We focus here on the case where $Z$ is univariate and continuous. To avoid the curse of dimensionality under some structural assumptions, for the multivariate case one can apply single index Fréchet regression (Bhattacharjee and Müller, 2023), employing arguments that are similar to those in Zhou and Müller (2024a) for the case of multivariate predictors in conformal inference. For the univariate case we use nonparametric regression with argument $Z$ to implement conditional profiles, employing local linear estimators (Fan and Gijbels, 1992) for $F_{\omega_{1},z}^{X}(u)$, $F_{\omega_{2},z}^{Y}(v)$, and $F_{\omega_{1},\omega_{2},z}^{XY}(u,v)$. This yields
@@
×$\quad\quad\quad\quad\times\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right),\hskip 369.88582pt$
where $\mathrm{K}(\cdot)$ is a symmetric and Lipschitz continuous density kernel on $[-1,1]$ and $h_{n}$ is a sequence of bandwidths.
@@
$\mathcal{F}_{1}$ =$=\,\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t\}}:\omega_{1}\in\mathcal{M}_{1}\,\,\,t\in{\mathbb{R}}^{+}\},$
$\mathcal{F}_{2}$ =$=\,\{f(y)=\mathds{1}_{\{d(\omega_{2},y)\leq t\}}:\omega_{2}\in\mathcal{M}_{2},\,\,\,t\in{\mathbb{R}}^{+}\},$
@@
For every $\omega_{1}\in\mathcal{M}_{1}$, $\omega_{2}\in\mathcal{M}_{2}$ and $u,v\in{\mathbb{R}}^{+}$, $F^{XY}_{\omega_{1},\omega_{2},z}(u,v)$ is second order differentiable and has bounded second order derivatives with respect to $z$. Moreover, there exists a constant $L^{\prime}$ such that $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1}^{\prime},\omega_{2},z}(u,v)|\leq L^{\prime}d(\omega_{1},\omega_{1}^{\prime})$, $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2}^{\prime},z}(u,v)|\leq L^{\prime}d(\omega_{2},\omega_{2}^{\prime})$ for all $z\in\mathcal{X},u,v\in{\mathbb{R}}^{+}$ and $\omega_{1},\omega_{1}^{\prime},\in\mathcal{M}_{1}$, $\omega_{2},\omega_{2}^{\prime},\in\mathcal{M}_{2}$.
@@
Under assumptions 1 and 2, if $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$ for some $v>0$ and $h_{n}\asymp(n/\log n)^{-1/5}$, then:
[…]
The first two statements of Proposition 4 follow directly from Theorem 3 in Zhou and Müller (2024a). Under the assumption $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$, Lemma A.1 in Einmahl and Mason (2000) and Lemma S.4 in Zhou and Müller (2024b) imply that the function class $\mathcal{F}_{12}=\left\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t_{1}\}}\mathds{1}_{\{d(\omega_{2},x)\leq t_{2}\}}:\omega_{1}\in\mathcal{M}_{1},\,\omega_{2}\in\mathcal{M}_{1},\,t_{1},t_{2}\in\mathbb{R}\right\}$ has a polynomial covering number. The third statement is a consequence of Lemma 1 and Theorem 3 in Zhou and Müller (2024a).
@@
[…]
@@
where $N$ denotes the total number of permutations. The randomized test for testing (11) at significance level $\alpha$ is then defined as $\psi_{\alpha}^{c}:=\mathbb{P}(p_{c}\leq\alpha)$.
Denoting by $\Gamma(z)$ the cumulative distribution function the limiting distribution in Theorem 6, its corresponding permutation-based estimator is
@@
Under the same assumptions as in Theorem 6, $|\hat{\Gamma}_{N}(z)-\Gamma(z)|=o_{P}(1)$ holds for every continuity point $z$ of $\Gamma(z)$, and $\mathbb{P}(\psi_{\alpha}^{c})\leq\alpha$.

### v3

@@ 3.1 Testing independence for random objects
We are interested in testing
$\mathcal{H}_{0}:X\perp\!\!\!\perp Y\quad\text{versus}\quad\mathcal{H}_{1}:X\not\!\perp\!\!\!\perp Y$
using random samples $\mathcal{D}_{\mathsf{XY}}:=\{(X_{i},Y_{i}):1\leq i\leq n\}$ drawn from the joint distribution of the random objects $(X,Y)$. The profile association $\Delta(X,Y)$ in (4) quantifies the strength of the relationship between random objects $X$ and $Y$ and provides the foundation for the development of a test for independence.
As discussed in Section 2, for scalar random variables $X$ and $Y$ with joint CDF $F$, Hoeffding (1948) introduced the functional $\Delta(F)=\iint\{F(x,y)-F(x,\infty)\,F(\infty,y)\}^{2}\,{\rm d}F(x,y)$ as a measure of dependence. This functional can be expressed in terms of indicator functions. Specifically, define $\psi(x_{1},x_{2},x_{3})=\mathds{1}_{\{x_{1}-x_{2}\geq 0\}}-\mathds{1}_{\{x_{1}-x_{3}\geq 0\}},$ where the indicator function $\mathds{1}_{A}=1$ if the event $A$ is true and $\mathds{1}_{A}=0$ otherwise. Setting $\phi(x_{1},y_{1};\ldots,x_{5},y_{5})=0.25\times\psi(x_{1},x_{2},x_{3})\psi(x_{1},x_{4},x_{5})\psi(y_{1},y_{2},y_{3})\psi(y_{1},y_{4},y_{5}),$ one can write
$\Delta(F)=\int\cdots\int\phi(x_{1},y_{1};\ldots,x_{5},y_{5})\,{\rm d}F(x_{1},y_{1})\cdots{\rm d}F(x_{5},y_{5}).$
When $X$ and $Y$ are random objects, we define the scalar random variables $U_{\omega_{1}}=d_{1}(X,\omega_{1})$ and $V_{\omega_{2}}=d_{2}(Y,\omega_{2})$, and denote their joint CDF by $\mathfrak{F}_{\omega_{1},\omega_{2}}$ for $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$. Using these notations, the profile association defined in (4) can be written as $\Delta(X,Y)=\mathbb{E}_{X^{\prime},Y^{\prime}}\left[\Delta\left(\mathfrak{F}_{X^{\prime},Y^{\prime}}\right)\right].$
Inspired by this, consider the functions $\psi_{1}:\mathcal{M}_{1}^{\otimes 4}\mapsto\{-1,0,1\}$ defined as
@@
$\quad\times\psi(y_{1},y_{2},y_{3},y_{4})\psi(y_{1},y_{2},y_{5},y_{6}).$
In what follows, we use the notation $\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{6},Y_{6})\}$ to denote six independent copies of $(X,Y)$. Following the previous discussion, we have
@@
Since the function $h$ is not symmetric, we first symmetrize it as follows,
@@
where $\mathcal{P}_{6}^{6}(\bm{\alpha})$ denotes the set of all permutations $\{\alpha_{1},\ldots,\alpha_{6}\}$ of $\{1,2,\ldots,6\}$. Given random samples $\mathcal{D}_{\mathsf{XY}}$, we employ a U-statistic as an estimator of $\Delta(X,Y)$, defined as
$D_{n}(\mathcal{D}_{\mathsf{XY}})=\binom{n}{6}^{-1}\sum_{\mathcal{C}_{6}^{n}(\bm{\alpha})}\mathfrak{h}\left((X_{\alpha_{1}},Y_{\alpha_{1}}),\ldots,(X_{\alpha_{6}},Y_{\alpha_{6}})\right),$ (5)
where $\sum_{\mathcal{C}_{6}^{n}(\bm{\alpha})}$ demotes summation over all $\binom{n}{6}$ combinations of $6$ distinct elements $\{\alpha_{1},\ldots,\alpha_{6}\}$ from $\{1,2,\ldots,n\}$.
We require some concepts related to the analysis of U-statistics, which were first introduced by Hoeffding (1948), with further details provided in Serfling (2009). For $r=1,2,\ldots,6$, let
@@
${n}(D_{n}(\mathcal{D}_{\mathsf{XY}})-\Delta(X,Y))\xrightarrow{d}{15}\sum_{j=1}^{\infty}\lambda_{j}(Z_{j}-1),$
@@
for all $(u,v)\in\mathcal{M}_{1}\times\mathcal{M}_{2}$ and $f\in\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\times\mathcal{M}_{2})$.
@@ 3.2 Permutation test and asymptotic results
Since the asymptotic distribution in Theorem 2 is a second-order Wiener chaos with unknown parameters, we use a permutation test to determine the critical values of the test statistic $D_{n}(\mathcal{D}_{\mathsf{XY}})$ when implementing the test. Permuting the data and recalculating the test statistic for each permutation is a standard approach to generate a reference distribution. In this subsection, we implement the proposed independence test using a permutation-based method and study its asymptotic properties, including consistency and power analysis. Recently, alternative approaches such as the random-lifter method (Gao et al., 2025) and spectrum-based methods (Wang et al., 2024) have been studied to handle independence testing in the presence of unknown limiting distributions. We include a discussion of the spectrum-based approach in the supplementary material.
@@
For testing independence, a common approach is to compute replicates of the test statistic by randomly permuting the $Y$ sample while keeping the order of the $X$ sample fixed (Székely et al., 2007). However, randomly permuting the entire $Y$ sample introduces dependence among the data pairs under the alternative $\mathcal{H}_{1}$ when $X$ and $Y$ are dependent, making it challenging to establish consistency of the critical value under $\mathcal{H}_{1}$. To address this issue, we adopt a half-permutation procedure introduced in Fan et al. (2024) that does not require a splitting technique with its possible loss of power. Adopting this approach for our setting requires a different proof technique to obtain consistency and power bounds for the proposed profile independence test, as the previous techniques are not applicable.
@@
We use $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$ to denote the set of all subsets of $\{1,2,\ldots,n\}$ of size $\lfloor n/2\rfloor$, where each subset consists of unrepeated random samples from $\{1,\ldots,n\}$. Consider a sequence of i.i.d. random permutations $\Pi_{1},\ldots,\Pi_{B}$ that are independent of $(X,Y)$ and uniformly distributed on $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$. We use $\Pi_{j}^{c}$ to denote the complement of $\Pi_{j}$ with respect to the set $\{1,2,\ldots,n\}$. Without loss of generality, we assume $n$ is even, so that $\Pi_{j}$ and $\Pi_{j}^{c}$ have the same cardinality $n/2$. For each $\Pi_{j}^{c}$, let $\{\sigma_{jk}\}_{k=1}^{N_{j}}$ be a sequence of random permutations on $\Pi_{j}^{c}$ that are independent of $(X,Y)$ and $\Pi_{j}$. Assume $\Pi_{j}=\{\imath_{1},\imath_{2},\ldots,\imath_{\lfloor n/2\rfloor}\}$ and $\sigma_{jk}\circ\Pi_{j}^{c}=\{\jmath_{1},\jmath_{2},\ldots,\jmath_{\lfloor n/2\rfloor}\}$, where $\sigma_{jk}\circ\Pi_{j}^{c}$ denotes the sequence of $\Pi_{j}^{c}$ under the permutation $\sigma_{jk}$. The permuted data corresponding to $(\Pi_{j},\sigma_{jk})$ are
@@
$\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}):=\{(X_{\imath_{1}},Y_{\jmath_{1}}),(X_{\imath_{2}},Y_{\jmath_{2}}),\ldots,(X_{\imath_{\lfloor n/2\rfloor}},Y_{\jmath_{\lfloor n/2\rfloor}})\},$
@@
and the test statistic based on permuted data $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ is
@@
$D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})):=\binom{\lfloor n/2\rfloor}{6}^{-1}\sum_{\mathcal{C}_{6}^{\lfloor n/2\rfloor}(\bm{\alpha})}h((X_{\imath_{\alpha_{1}}},Y_{\jmath_{\alpha_{1}}}),\ldots,(X_{\imath_{\alpha_{6}}},Y_{\jmath_{\alpha_{6}}})).$
@@
$\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}}):=\mathds{1}\left(1+\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\{n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})\leq\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}}\leq(1+N)\alpha\right),$ (8)
where the only randomness in $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ comes from the random permutations $(\Pi_{j},\sigma_{jk})$ and $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})=1$ indicates that the test rejects $H_{0}$ at level $\alpha$.
Under the null hypothesis, $X$ and $Y$ are independent, and the elements in the data sets $\mathcal{D}_{\mathsf{XY}}$ and $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ are exchangeable. Consequently, $\{\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}_{j,k}$ and $n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})$ have the same asymptotic distribution given by Theorem 2. Therefore, the probability on the right-hand side of equation (8) equals $\alpha$ for all distributions of $(X,Y)$ under the null hypothesis, which implies that $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ is a valid size-$\alpha$ test. For the consistency under the alternative $\mathcal{H}_{1}$, denote by $H(z)$ the cumulative distribution function of the asymptotic null distribution in Theorem 2. We employ the estimator
@@
for $H(z)$ based on the permuted data. Theorem 3 below establishes the consistency of this estimator, which ensures that the size of the proposed profile independence test is correctly controlled under the alternative.
@@ 4 Conditional independence test
Conditional independence testing is a statistical tool used to determine whether two variables are independent when conditioning on a third variable. It is used in a wide range of statistical applications where one studies complex relationships among variables, including causal inference (Zhang et al., 2011, 2017; Peters et al., 2017), graphical models (Koller and Friedman, 2009; Williams and Mulder, 2020) and sufficient dimension reduction (Cook, 2004; Chen et al., 2015; Li, 2018). Many nonparametric conditional independence tests have been studied based on conditional density functions (Huang, 2010; Su and White, 2008; Banerjee et al., 2024), distribution functions (Linton and Gozalo, 1996; Zhou et al., 2020) and characteristic functions (Su and White, 2007; Wang et al., 2015; Wang and Hong, 2018). These methods are all restricted to the case where data reside in a vector space. To our knowledge, conditional independence tests for random elements in general metric spaces that do not support linear operations have remained unexplored.
We consider a setting where $X\in\mathcal{M}_{1}$ and $Y\in\mathcal{M}_{2}$ are random objects, while the conditional variable $Z$ is Euclidean, specifically $Z\in\mathcal{Z}$, where $\mathcal{Z}$ is a compact subset of ${\mathbb{R}}^{p}$. Our goal is to test the following hypothesis:
$\mathcal{H}_{0,c}:X\mid Z\perp\!\!\!\perp Y\mid Z\quad\text{versus}\quad\mathcal{H}_{1,c}:X\mid Z\not\!\perp\!\!\!\perp Y\mid Z,$ (11)
@@
$F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v)=\mathbb{P}_{X,Y,Z}\left(d(X,X^{\prime})\leq u,\,d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right),$
$F_{X^{\prime},Z^{\prime}}^{X}(u)=\mathbb{P}_{X,Z}\left(d(X,X^{\prime})\leq u\mid Z=Z^{\prime}\right),\,F_{Y^{\prime},Z^{\prime}}^{Y}(v)=\mathbb{P}_{Y,Z}\left(d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right).$
@@
The definition (12) places no restrictions on the distribution of $Z$; hence, it remains valid whether $Z$ is continuous, discrete, or of mixed type. We focus on the case where $Z$ is univariate and continuous in the estimation procedure, as this setting is both commonly encountered in practice and theoretically nontrivial, particularly due to the challenges involved in establishing uniform convergence for local linear smoothing with object-valued data. When $Z$ is discrete (e.g., categorical), the estimation becomes significantly simpler: one can compute the association measure within each category of $Z$ and aggregate the results, avoiding the need for smoothing altogether. To avoid the curse of dimensionality under some structural assumptions, for the multivariate case one can apply single index Fréchet regression (Bhattacharjee and Müller, 2023), employing arguments that are similar to those in Zhou and Müller (2025) for the case of multivariate predictors in conformal inference. For the case of a univatiate continuous predictor $Z$ we use nonparametric regression with argument $Z$ to implement conditional profiles, employing local linear estimators (Fan and Gijbels, 1992) for $F_{\omega_{1},z}^{X}(u)$, $F_{\omega_{2},z}^{Y}(v)$, and $F_{\omega_{1},\omega_{2},z}^{XY}(u,v)$. This yields
@@
$\quad\quad\quad\quad\times\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right),\hskip 369.88582pt$
where $\mathrm{K}(\cdot)$ is a symmetric and Lipschitz continuous density kernel on [1,1][-1,1] and $h_{n}$ is a sequence of bandwidths.
@@
$\mathcal{F}_{1}$ $=\,\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t\}}:\omega_{1}\in\mathcal{M}_{1}\,\,\,t\in{\mathbb{R}}^{+}\},$
$\mathcal{F}_{2}$ $=\,\{f(y)=\mathds{1}_{\{d(\omega_{2},y)\leq t\}}:\omega_{2}\in\mathcal{M}_{2},\,\,\,t\in{\mathbb{R}}^{+}\},$
@@
For every $\omega_{1}\in\mathcal{M}_{1}$, $\omega_{2}\in\mathcal{M}_{2}$ and $u,v\in{\mathbb{R}}^{+}$, $F^{XY}_{\omega_{1},\omega_{2},z}(u,v)$ is second order differentiable and has bounded second order derivatives with respect to $z$. Moreover, there exists a constant $L^{\prime}$ such that $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1}^{\prime},\omega_{2},z}(u,v)|\leq L^{\prime}d_{1}(\omega_{1},\omega_{1}^{\prime})$, $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2}^{\prime},z}(u,v)|\leq L^{\prime}d_{2}(\omega_{2},\omega_{2}^{\prime})$ for all $z\in\mathcal{X},u,v\in{\mathbb{R}}^{+}$ and $\omega_{1},\omega_{1}^{\prime},\in\mathcal{M}_{1}$, $\omega_{2},\omega_{2}^{\prime},\in\mathcal{M}_{2}$.
@@
Under assumptions 1 and 4, if $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$ for some $v>0$ and $h_{n}\asymp(n/\log n)^{-1/5}$,
$\sup_{\omega_{1}\in\mathcal{M}_{1}}\sup_{z\in\mathcal{Z}}\sup_{u\geq 0}\left|\hat{F}^{X}_{\omega_{1},z}(u)-F^{X}_{\omega_{1},z}(u)\right|=O\left(\left(\frac{\log n}{n}\right)^{\frac{2}{5}}\right)\text{ a.s.},$
$\sup_{\omega_{2}\in\mathcal{M}_{2}}\sup_{z\in\mathcal{Z}}\sup_{v\geq 0}\left|\hat{F}^{Y}_{\omega_{2},z}(v)-F^{Y}_{\omega_{2},z}(v)\right|=O\left(\left(\frac{\log n}{n}\right)^{\frac{2}{5}}\right)\text{ a.s.},$
$\sup_{\omega_{1}\in\mathcal{M}_{1}}\sup_{\omega_{2}\in\mathcal{M}_{2}}\sup_{z\in\mathcal{Z}}\sup_{u,v\geq 0}\left|\hat{F}^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2},z}(u,v)\right|=O\left(\left(\frac{\log n}{n}\right)^{\frac{2}{5}}\right)\text{ a.s.}.$
@@
$\sup_{z\in\mathcal{Z}}|\hat{\Delta}_{XY}(z)-{\Delta}_{XY}(z)|=O\left(h^{2}+\sqrt{\frac{\log n}{nh_{n}}}\right)\text{ a.s.}.$
@@
where $N$ denotes the total number of permutations. The randomized test for testing (11) at significance level $\alpha$ is then defined as $\psi_{\alpha}^{c}:=\mathds{1}(p_{c}\leq\alpha)$.
Denoting by $\Gamma(z)$ the cumulative distribution function of the limiting distribution in Theorem 6, its corresponding permutation-based estimator is
@@
Under the same assumptions as in Theorem 6, $|\hat{\Gamma}_{N}(z)-\Gamma(z)|=o_{P}(1)$ holds for every continuity point $z$ of $\Gamma(z)$.